OpenStudy (anonymous):

Solve the compound inequality, graph the solution. -2 < (greater or less than) 2x -4 <4

OpenStudy (anonymous):

what do you suggest we do first

OpenStudy (anonymous):

\[-2\le2x-4<4\] how do we get x by itself

OpenStudy (anonymous):

subtract it from both sides?

OpenStudy (anonymous):

noo dont bother with the x right now look at whats next to it

OpenStudy (anonymous):

(-4)

OpenStudy (anonymous):

You asked how to get x by itself

OpenStudy (anonymous):

yes but we dont go straight to x we need to get rid of everything around it then deal with the coefficient

OpenStudy (anonymous):

So add four to each side

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[2\le2x <8\]

OpenStudy (anonymous):

So now its -2 < 2x +8

OpenStudy (anonymous):

Nevermind, i wrote it down wrong

OpenStudy (anonymous):

Now you add 2 to both sides?

OpenStudy (anonymous):

now to isolate x we do the reciprocal of the coefficient

OpenStudy (anonymous):

whats the reciprocal of multiplying?

OpenStudy (anonymous):

dividing

OpenStudy (anonymous):

yes so we divide everything by 2

OpenStudy (anonymous):

\[1\le x <8\]

OpenStudy (anonymous):

so 1 < x < 4

OpenStudy (anonymous):

I divided the 8 by two

OpenStudy (anonymous):

\[1\le x <4\]

OpenStudy (anonymous):

yes i forgot to divide the 8 when i wrote it

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now subrtract the 1?

OpenStudy (anonymous):

no that would be the answer

OpenStudy (anonymous):

The answer is graphed I

OpenStudy (anonymous):

http://gyazo.com/24f89d4464beb58f0c6d792c63e2f405

OpenStudy (anonymous):

yes it's graphed with an inequality which one has our inequality \[1\le x <4\]

OpenStudy (anonymous):

\[\le = closed <=open\]